OpenStudy (inhaleof):

Solve the inequality -5x + 8 less than or greater to x - 4 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (inhaleof):

-5x+8<=x-4

OpenStudy (jdoe0001):

so.... if you solve for "x"... what would you get?

OpenStudy (anonymous):

a headache

OpenStudy (radar):

Say what you mean, mean what you say.\[-5x + 8 <> x - 4 \] or do you mean to say\[-5x + 8 \le x - 4\]

OpenStudy (radar):

\[- 5x + 8 \neq x - 4\] if it is what you "state" in the problem.

OpenStudy (radar):

\[-5x+8 \le x - 4\]Use Algebra: \[-6x \le -12\]\[x \ge 2\] Now graph it.

OpenStudy (radar):

This is assuming that your second post is what you mean.